OpenStudy (anonymous):

what is the slope of a line parallel to 8/9?

OpenStudy (anonymous):

\[\frac{8}{9}\] is a number not a line

OpenStudy (anonymous):

\(y=\frac{8}{9}x+22\) is a line also \(y=\frac{8}{9}\) is a line, but just \(\frac{8}{9}\) is not a line